OpenStudy (anonymous):

If f(x) = 1/3 x − 2 and g(x) = −2x − 3, find (f • g)(6) 15 −15 0 −17

hero (hero):

(f • g)(x) = f(x) • g(x) = (x-2)/3 • (-2x-3) (f • g)(6) = f(6) • g(6) = (6-2)/3 • (-2(6) -3) = 3/3 • -12 - 3 = 1 • -15 = -15

OpenStudy (anonymous):

thanx